# Constraining Atmospheric River Uncertainty Using Instantaneous Poleward Latent Heat Transport

### v1

@@ Abstract
Atmospheric rivers (ARs) are extreme weather events that play a crucial role in the global hydrological cycle. As a key mechanism of latent heat transport (LHT), they help maintain energy balance in the climate system. While an AR is characterized by a long, narrow corridor of water vapor associated with a low-level jet stream, there is no unambiguous definition of an AR grounded in geophysical fluid dynamics. Therefore, AR identification is currently performed by a large array of expert-defined, threshold-based algorithms. The variety of algorithms introduces uncertainty in the estimated contribution of ARs to LHT. We calculate the instantaneous eddy LHT from moist, poleward anomalies. Based on the dynamics of the large-scale atmospheric circulation, this quantity is a physics-based upper bound that constrains AR projections from the variety of detection algorithms. We quantify the contribution of ARs to transient eddies, stationary eddies, and transient-stationary eddy interactions, and we show the relative contribution of ARs and other processes, such as dry, equatorward transport. We use this upper bound to quantify ARs’ frequency, intensity, and temporal variability. In the historical climate, we find that ARs transport $\sim$2.21 PW at the latitude of peak meridional transport in Northern Hemisphere winter, with approximately 0.47 PW of temporal variability. By the end of the century in a future climate projection, at this latitude, we find that AR-induced LHT will increase by 0.5 PW and the corresponding temporal variability will increase by 0.14 PW.
@@ 1 Introduction
In the climate system, incoming solar radiation and outgoing longwave radiation are not equal at each latitude, and poleward energy transport is necessary to maintain energy balance (Peixoto et al., 1992). Atmospheric poleward energy transport can occur in the form of sensible heat, latent heat, and geopotential. Atmospheric rivers (ARs) are long, narrow filaments of moisture transport that are a mechanism of poleward latent heat transport (LHT) (Zhu and Newell, 1998). These extreme events can cause record-breaking floods (DeFlorio et al., 2024; Corringham et al., 2019). For some regions, they play a crucial role as “drought-busters,” in which they are responsible for the vast majority of rainfall (Dettinger, 2013). Despite their significance, these events do not have an unambiguous definition grounded in geophysical fluid dynamics, and various AR detection algorithms have been presented in the Atmospheric River Tracking Method Intercomparison Project (ARTMIP) (Rutz et al., 2019; Shields et al., 2018). At a single time step, the AR detection uncertainty is visualized as a function of ARTMIP consensus in Figure 1 of Mahesh et al. (2024). Existing studies in ARTMIP have quantified the detection uncertainty of ARs in different reanalysis datasets (Collow et al., 2022), climate models (Shields et al., 2023; O’Brien et al., 2022), and climate scenarios (Shields et al., 2021; Rush et al., 2025). The uncertainty from the choice of detection algorithm leads to significant changes in AR-induced Integrated Vapor Transport (IVT), with an uncertainty ranging from 400 to 800 kg/m/s (Figure 2 of Shields et al. (2023)). This has led to uncertainties in the past, present, and future behavior of ARs in the midlatitudes (Lora et al., 2020; Zhou et al., 2021) and polar regions (Shields et al., 2022; Maclennan et al., 2025; Thaker et al., 2025).
@@ 4.1 A Method to Calculate the Transport from Moist, Poleward Anomalies
4.2 Comparing Eddy LHT from ARTMIP Algorithms and Moist, Poleward Anomalies
@@
where the subscripts M denotes moist, and P denotes poleward. The reference quantity for AR-induced LHT is the moist, poleward component of TEs, SEs, and MTs. To put this quantity in the broader context of poleward heat transport, we present a flowchart in Figure 4. This chart shows the ways to decompose poleward heat transport to derive AR reference LHT in Equation 16. We show that this quantity serves as an upper bound to AR-induced transport across all ARTMIP algorithms (Figure 2c). In DJF (Figure 2), the Northern Hemisphere latitude of peak AR Reference LHT is 34 N, which has an AR Reference LHT of 2.21 PW, with a temporal standard deviation of 0.47 PW. The Southern Hemisphere latitude of peak AR Reference LHT is 40.5 S, with 2.68 PW of transport, with a temporal standard deviation of 0.59 PW. In JJA (Appendix Figure A1), the Northern Hemisphere latitude of peak AR Reference LHT is 43.5 N, which has an AR Reference LHT of 1.75 PW with a temporal standard deviation of 0.51 PW. The Southern Hemisphere latitude of peak AR Reference LHT is 39.5 S, with 2.61 PW of transport, with a temporal standard deviation of 0.49 PW
@@ 6.1 Latent Heat Transport
Figure 7 shows the DJF and JJA changes in the AR reference LHT, comparing the 2081-2100 period under the SSP5-8.5 scenario to the 1996-2014 historical scenario. In DJF, the AR Reference LHT increases by 0.5 PW at 34 N and by 0.43 PW at 40.5 S. In JJA, the AR Reference LHT increases by 0.25 PW at 43.5 N and 0.56 PW at 39.5 S. These latitudes are the latitudes of peak transport in the Northern and Southern Hemispheres discussed in Section 3a. The changes from the moist, poleward anomalies are greater than those from the other anomalies. For the total component, the climate changes from the mixed terms are also 0, so the future climate change in the total component comes from TEs and SEs. In the Northern Hemisphere summer, the increase in TE and AR reference LHT is muted due to the increase in stationary eddy transport (Donohoe et al., 2020).
@@
The historical variability of AR reference LHT is similar between MERRA2 (Figure 3) and the MRI earth system model (Figure 8). Under future climate conditions, the temporal variability of AR reference LHT increases (Figure 8). This increase is approximately 0.1-0.2 PW in the extratropics during both seasons. We discuss possible relations between the increase in variability and the meridional moisture gradient in the Discussion Section.
@@ 6.2 Frequency
Under climate change, a greater proportion of moist, poleward anomalies are classified as ARs (Figure 10 and Appendix Figure A7). In both DJF and JJA, the conditional probability of an AR, given a moist, poleward anomaly increases significantly in the future climate, particularly for the GuanWaliser_v2, Lora_v2, and Mundhenk algorithms. However, there is still spatial variability, such as a detected decrease in AR frequency over the Pacific Ocean in DJF (though not JJA) in the ARCONNECT_v2 algorithm. Under future climate change, a greater proportion of moist, poleward anomalies reach the threshold to be classified as ARs.
@@ 7 Discussion and Conclusions
In this manuscript, we focus on the relationship between ARs and LHT because we assess the role of ARs in the global water cycle. However, ARs have also been associated with sensible heat transport (Shields et al., 2019) and surface heat extremes (Scholz and Lora, 2024; Raymond et al., 2024). Future research is necessary to assess the role of ARs and moist, poleward anomalies in dry static energy transport. This could improve understanding of the role of ARs in column-integrated moist static energy tendencies, high heat-humidity events in the midlatitudes, and polar climate change. In particular, moist and dry heat transport changes have distinct seasonality, with important implications for Arctic warming and sea ice loss (Hahn et al., 2023). Additionally, our manuscript quantifies the relative contributions of moist, poleward anomalies; dry, equatorward anomalies; and different-signed anomalies in LHT. A future analysis could quantify the analogous contributions from warm, poleward anomalies and cold, equatorward anomalies in TE dry static energy transport.
@@ Acknowledgements.
ARTMIP is a grassroots community effort that includes a collection of international researchers from universities, laboratories, and agencies. Co-chairs and committee members include Jonathan Rutz, Christine Shields, L. Ruby Leung, F. Martin Ralph, Michael Wehner, Ashley Payne, Travis O’Brien, and Allison Collow. Details on the catalog developers can be found on the ARTMIP website (http://www.cgd.ucar.edu/projects/artmip, last access: 12 April 2025). ARTMIP has received support from the U.S. Department of Energy Office of Science Biological and Environmental Research (BER), as part of the Regional and Global Climate Modeling program, and the Center for Western Weather and Water Extremes (CW3E) at Scripps Institute for Oceanography at the University of California San Diego. This research has been supported by the director, Office of Science, Office of Biological and Environmental Research of the U.S. Department of Energy, as part of the Regional and Global Model Analysis program area, within the Earth and Environmental Systems Modeling Program (contract no. DE-AC02-05CH11231), and used resources of the National Energy Research Scientific Computing Center (NERSC), also supported by the Office of Science of the U.S. Department of Energy (contract no. DE-AC02-05CH11231). This research was also supported in part by the Environmental Resilience Institute, funded by Indiana University’s Prepared for Environmental Change Grand Challenge initiative, and in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute. [A] We use the following AR detection algorithms: ar-connect (Shearer et al., 2020), guan-waliser (Guan and Waliser, 2015), lora_v2 (Skinner et al., 2020; Lora et al., 2017), mattingly_v2 (Mattingly et al., 2018), mundhenk_v3 (Mundhenk et al., 2016), panlu (Pan and Lu, 2019), reid (Reid et al., 2020), rutz (Rutz et al., 2014), connect (Sellars et al., 2015), sail_v1 (Gavrikov et al., 2020), SCAFET (Nellikkattil et al., 2023), teca_bard_v1.0.1 (O'Brien et al., 2020), and tempest (McClenny et al., 2020; Ullrich et al., 2021). [B]
@@ Appendix B Decomposition of TE LHT into 3 sub-terms
Based on Equations 7 and  9,

### v2

@@ Abstract
Atmospheric rivers (ARs) are extreme weather events that play a crucial role in the global hydrological cycle. As a key mechanism of latent heat transport (LHT), they help maintain energy balance in the climate system. While an AR is characterized by a long, narrow corridor of water vapor associated with a low-level jet stream, there is no unambiguous definition of an AR grounded in geophysical fluid dynamics. AR identification is currently performed by a variety of threshold-based algorithms, which has introduced uncertainty in the estimated contribution of ARs to LHT. We calculate the instantaneous eddy LHT from moist, poleward anomalies. Based on the dynamics of the large-scale atmospheric circulation, this quantity is a physics-based upper bound that constrains AR projections from the variety of detection algorithms. We quantify the contribution of ARs to transient eddies, stationary eddies, and transient-stationary eddy interactions, and we show the relative contributions of ARs vs. other processes, such as dry, equatorward transport. We use this upper bound as a reference to quantify ARs’ frequency, intensity, and temporal variability. In the historical climate, the AR reference transport is $\sim$2.21 PW at the latitude of peak transport in Northern Hemisphere winter, with a temporal standard deviation of approximately 0.47 PW. In a future climate projection, at this latitude, AR-induced LHT will increase by 0.5 PW and the corresponding temporal variability will increase by 0.14 PW. The future change in the AR reference transport is larger than the changes associated with other types of atmospheric anomalies.
@@ 1 Introduction
In the climate system, incoming solar radiation and outgoing longwave radiation are not equal at each latitude, and poleward energy transport is necessary to maintain energy balance (Peixoto et al., 1992). Atmospheric poleward energy transport can occur in the form of sensible heat, latent heat, and geopotential. Atmospheric rivers (ARs) are long, narrow filaments of moisture transport that are a mechanism of poleward latent heat transport (LHT) (Zhu and Newell, 1998). These extreme events can cause record-breaking floods (DeFlorio et al., 2024; Corringham et al., 2019). For some regions, they play a crucial role as “drought-busters,” in which they are responsible for the vast majority of rainfall (Dettinger, 2013). Despite their significance, these events do not have an unambiguous definition grounded in geophysical fluid dynamics, and various AR detection algorithms have been presented in the Atmospheric River Tracking Method Intercomparison Project (ARTMIP) (Rutz et al., 2019; Shields et al., 2018). At a single time step, the AR detection uncertainty is visualized as a function of ARTMIP consensus in Figure 1 of Mahesh et al. (2024). Existing studies in ARTMIP have quantified the detection uncertainty of ARs in different reanalysis datasets (Collow et al., 2022), climate models (Shields et al., 2023; O’Brien et al., 2022), and climate scenarios (Shields et al., 2021; Rush et al., 2025). Depending on the detection algorithm used, estimates of AR-induced Integrated Vapor Transport (IVT) vary widely, ranging from 400 to 800 kg/m/s (Figure 2 of Shields et al. (2023)). This has led to significant uncertainties in the past, present, and future behavior of ARs in the midlatitudes (Lora et al., 2020; Zhou et al., 2021) and polar regions (Shields et al., 2022; Maclennan et al., 2025; Thaker et al., 2025).
@@ 4.1 A Method to Calculate the Transport from Moist, Poleward Anomalies
4.2 Comparing Eddy LHT from ARTMIP Algorithms and from Moist, Poleward Anomalies
@@
where the subscript M denotes moist, and P denotes poleward. The reference quantity for AR-induced LHT is the moist, poleward component of TEs, SEs, and MTs. To put this quantity in the broader context of poleward heat transport, we present a flowchart in Figure 4. This chart shows the ways to decompose poleward heat transport to derive AR reference LHT in Equation 16. We show that this quantity serves as an upper bound to AR-induced transport across all ARTMIP algorithms (Figure 2c). In DJF (Figure 2), the Northern Hemisphere latitude of peak AR Reference LHT is 34 N, which has an AR Reference LHT of 2.21 PW, with a temporal standard deviation of 0.47 PW. The Southern Hemisphere latitude of peak AR Reference LHT is 40.5 S, with 2.68 PW of transport, with a temporal standard deviation of 0.59 PW. In JJA (Appendix Figure A1), the Northern Hemisphere latitude of peak AR Reference LHT is 43.5 N, which has an AR Reference LHT of 1.75 PW with a temporal standard deviation of 0.51 PW. The Southern Hemisphere latitude of peak AR Reference LHT is 39.5 S, with 2.61 PW of transport, with a temporal standard deviation of 0.49 PW.
@@ 6.1 Latent Heat Transport
Figure 7 shows the DJF and JJA changes in the AR reference LHT, comparing the 2081-2100 period under the SSP5-8.5 scenario to the 1996-2014 historical scenario. In DJF, the AR Reference LHT increases by 0.5 PW at 34 N and by 0.43 PW at 40.5 S. In JJA, the AR Reference LHT increases by 0.25 PW at 43.5 N and 0.56 PW at 39.5 S. These latitudes are the latitudes of peak transport in the Northern and Southern Hemispheres discussed in Section 3a. By definition, the historical average of the MTs is 0, and the climate change from the MTs is 0. (The moist, poleward component of the historical and future MTs can be nonzero, however.) Therefore, the future climate change in the total component comes from TEs and SEs. In the Northern Hemisphere summer, the increase in TE and AR reference LHT is muted due to the increase in stationary eddy transport (Donohoe et al., 2020). Across all latitudes and both DJF and JJA, the change in moist, poleward transport is greater than the change in other transport (Figure 7). This indicates that the AR Reference LHT is an important driver of future changes in eddy LHT.
@@
The historical variability of AR reference LHT is similar between MERRA2 (Figure 3) and the MRI earth system model (Figure 8). Under future climate conditions, the temporal variability of AR reference LHT increases (Figure 8). This increase is approximately 0.1-0.2 PW in the extratropics during both seasons. We discuss possible relations between the increase in variability and the meridional moisture gradient in the Discussion.
@@ 6.2 Frequency
Under climate change, a greater proportion of moist, poleward anomalies are classified as ARs (Figure 10 and Appendix Figure A7). In both DJF and JJA, the conditional probability of an AR, given a moist, poleward anomaly, increases significantly in the future climate, particularly for the GuanWaliser_v2, Lora_v2, and Mundhenk algorithms. However, there is still spatial variability, such as a detected decrease in AR frequency over the Pacific Ocean in DJF (though not JJA) in the ARCONNECT_v2 algorithm. Under future climate change, a greater proportion of moist, poleward anomalies reach the threshold to be classified as ARs.
@@ 7 Discussion and Conclusions
In this manuscript, we focus on the relationship between ARs and LHT because we assess the role of ARs in the global water cycle. However, ARs have also been associated with sensible heat transport (Shields et al., 2019) and surface heat extremes (Scholz and Lora, 2024; Raymond et al., 2024). Future research is necessary to assess the role of ARs and moist, poleward anomalies in dry static energy transport. This could improve understanding of the role of ARs in column-integrated moist static energy tendencies, high heat-humidity events in the midlatitudes, and polar climate change. In particular, moist and dry heat transport changes have distinct seasonality, with important implications for Arctic warming and sea ice loss (Hahn et al., 2023). Additionally, our manuscript quantifies the relative contributions of moist, poleward anomalies; dry, equatorward anomalies; and different-signed anomalies in LHT. A future direction of research is to quantify the analogous contributions from warm, poleward anomalies and cold, equatorward anomalies in TE dry static energy transport. This would enable a deeper understanding of the contrast between poleward and equatorward transport in a variety of climate states (Caballero and Hanley, 2012; Smith et al., 2021).
@@ Acknowledgements.
ARTMIP is a grassroots community effort that includes a collection of international researchers from universities, laboratories, and agencies. Co-chairs and committee members include Jonathan Rutz, Christine Shields, L. Ruby Leung, F. Martin Ralph, Michael Wehner, Ashley Payne, Travis O’Brien, and Allison Collow. Details on the catalog developers can be found on the ARTMIP website (http://www.cgd.ucar.edu/projects/artmip, last access: 12 April 2025). ARTMIP has received support from the U.S. Department of Energy Office of Science Biological and Environmental Research (BER), as part of the Regional and Global Climate Modeling program, and the Center for Western Weather and Water Extremes (CW3E) at Scripps Institute for Oceanography at the University of California San Diego. This research has been supported by the director, Office of Science, Office of Biological and Environmental Research of the U.S. Department of Energy, as part of the Regional and Global Model Analysis program area, within the Earth and Environmental Systems Modeling Program (contract no. DE-AC02-05CH11231), and used resources of the National Energy Research Scientific Computing Center (NERSC), also supported by the Office of Science of the U.S. Department of Energy (contract no. DE-AC02-05CH11231). This research was also supported in part by the Environmental Resilience Institute, funded by Indiana University’s Prepared for Environmental Change Grand Challenge initiative, and in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute. [A] We use the following AR detection algorithms: ar-connect (Shearer et al., 2020), guan-waliser (Guan and Waliser, 2015), lora_v2 (Skinner et al., 2020; Lora et al., 2017), mattingly_v2 (Mattingly et al., 2018), mundhenk_v3 (Mundhenk et al., 2016), panlu (Pan and Lu, 2019), reid (Reid et al., 2020), rutz (Rutz et al., 2014), connect (Sellars et al., 2015), sail_v1 (Gavrikov et al., 2020), SCAFET (Nellikkattil et al., 2023), teca_bard_v1.0.1 (O'Brien et al., 2020), and tempest (McClenny et al., 2020; Ullrich et al., 2021).
Figure 11: Decomposing Eddy Latent Heat Transport. Same as Figure 2, but for JJA.
Figure 12: Absolute Contributions of Moist Poleward, Dry Equatorward, and Different Sign $v$ and $q$ anomalies to transient eddy transports. In MERRA2 (1995-2014), the contributions of moist, poleward; dry, equatorward; and different-sign anomalies are shown to transient eddy LHT. Contributions are shown for DJF (left) and JJA (right). See Figure 5 for the relative contributions.
Figure 13: JJA Temporal Distribution of total LHT, transient eddy LHT, and AR Upper Bound LHT. Same as Figure 6 but for JJA.
Figure 14: Histograms of vIVT (top row) and AR reference LHT and TE LHT (bottom row). Visualization of distributions corresponding to the skewness and kurtosis values shown in Figure 6). Top row shows vIVT, which is a zonally resolved quantity. Bottom row shows AR reference LHT and TE LHT, which are zonal mean quantities. Skewness and kurtosis values for each distribution are printed alongside them.
Figure 15: Decomposing Eddy Latent Heat Transport in MRI-ESM2-0. The sum of transient eddies, stationary eddies, and mixed terms are shown for DJF and JJA in MRI-ESM2-0 historical (1995-2014) and SSP5-8.5 (2081-2100).
Figure 16: Climate Changes in Frequency of ARs and Moist, Poleward Anomalies. Same as Figure 16 but for JJA.
Figure 17: Climate Changes in Conditional Probability of Atmospheric Rivers, Given Moist, Poleward Anomalies. Same as Figure 10 but for JJA.
[B]
@@ Appendix B Decomposition of TE LHT into 3 sub-terms
First, by definition,
[…]
$1-H(q^{\prime*})=H(-q^{\prime*})$
Therefore,
$1-H(\tilde{v}^{\prime*})H(q^{\prime*})$ $=(1-H(\tilde{v}^{\prime*}))H(q^{\prime*})+H(\tilde{v}^{\prime*})(1-H(q^{\prime*}))+(1-H(\tilde{v}^{\prime*}))(1-H(q^{\prime*}))$
$=H(-\tilde{v}^{\prime*})H(q^{\prime*})+H(\tilde{v}^{\prime*})H(-q^{\prime*})+H(-\tilde{v}^{\prime*})H(-q^{\prime*})$ (20)
Based on Equations 7,  9, and Equation 20